OpenStudy (anonymous):

given the function (shown down below) describe the transformation, find the domain, range, and graph

OpenStudy (anonymous):

\[y= \sqrt{x-2}+3\]

OpenStudy (anonymous):

the parent function is \[\large y = \sqrt{x} \] this parameter "a" translates the graph left or right depending on the value of a: \[\large y = \sqrt{x-a} \] this parameter "k" translates the graph up or down depending on the value of k: \[\large y = \sqrt{x} +k\] so with that info, what do you think the graph will look like if your equation is: \[\large y = \sqrt{x-2}+3 \]